can you tell me how to graph the points of y=3x-1. I get confused on the -1 part. is that negative 1 or minus 1. thank you
Answer by mmm4444bot(95)

Hello There:
As far as the -1 goes, there are two ways of looking at it. It is subtracting a positive 1, or you can say that it is adding a negative 1.
y = 3*x - (+1)
y = 3*x + (-1)
Both interpretations mean the same thing.
I tend to see it as subtracting a positive 1.
To graph the solutions to this equation (a line), we need to find the coordinants of two points that satisfy the equation. Then it's just a matter of plotting these two points and drawing a straight line through them.
Two easy points to find are the x- and y-intercepts. The x-intercept occurs when y is zero, and the y-intercept occurs when x is zero.
Let's find the y-intercept first. Set x equal to zero.
y = 3*(0) - 1
y = -1
The y-intercept is (0, -1)
Now we find the x-intercept. Set y equal to zero.
0 = 3*x - 1
Add 1 to both sides.
0 + 1 = 3*x + 1 - 1
3*x = 1
Divide both sides by 3.
(3/3)*x = 1/3
x = 1/3
The x-intercept is (1/3, 0).
If you find that these two points are too close together to draw an accurate line for your graph, then you can just pick some values for x and substitute them into the equation to get other solutions.
For example, let's choose 10 for x.
y = 3*(10) - 1
y = 30 - 1
y = 29
So, the point (10, 29) is also a point on the line.
~ Mark